### v1

# Enhancing Large Language Model-based Speech Recognition by Contextualization for Rare and Ambiguous Words

Abstract
We develop a large language model (LLM) based automatic speech recognition (ASR) system that can be contextualized by providing keywords as prior information in text prompts. We adopt decoder-only architecture and use our in-house LLM, PLaMo-100B, pre-trained from scratch using datasets dominated by Japanese and English texts as the decoder. We adopt a pre-trained Whisper encoder as an audio encoder, and the audio embeddings from the audio encoder are projected to the text embedding space by an adapter layer and concatenated with text embeddings converted from text prompts to form inputs to the decoder. By providing keywords as prior information in the text prompts, we can contextualize our LLM-based ASR system without modifying the model architecture to transcribe ambiguous words in the input audio accurately. Experimental results demonstrate that providing keywords to the decoder can significantly improve the recognition performance of rare and ambiguous words.
1 Introduction
In recent years, large language models (LLMs) have been attracting attention thanks to their very high performance in natural language processing (Devlin et al., 2019; Brown et al., 2020; Llama Team, AI @ Meta, 2024). The performance of LLMs can improve by increasing the amount of calculations, the amount of data, and the number of model parameters, as known as scaling law (Kaplan et al., 2020). Such pre-trained LLMs can solve different downstream tasks by contextualizing input texts, namely, prompts (Brown et al., 2020) without updating parameters. In addition, supervised fine-tuning can further improve the performance of LLMs for downstream tasks (Wei et al., 2022). However, supervised fine-tuning requires many computational resources to update the parameters of LLMs. Often, parameter efficient fine-tuning (Ding et al., 2022), where the subset of parameters is fixed, compromises the performance of downstream tasks and computational resources: the amount of memory.
There have also been many attempts to use voice input for decoder-only LLMs to have the multi-modal capability (Latif et al., 2023). In these models, outputs from audio encoders are discretized (Zhang et al., 2023; Rubenstein et al., 2023) like text tokens in LLMs or converted to token embeddings of LLMs by applying additional adapter modules (Shu et al., 2023; Wu et al., 2023; Fathullah et al., 2024), which are parameterized by shallow neural networks, and concatenated with text embeddings as inputs of LLMs. Thanks to the contextualization of LLMs by prompting, LLMs can perform not only automatic speech recognition (ASR) but also other tasks, such as speech translation and voice chat (Chu et al., 2024; Llama Team, AI @ Meta, 2024) by including instructions in the prompt texts to the LLMs.
Even to specialize ASR tasks, the LLM-based ASR systems can be contextualized by providing prior information on input speech to the prompts without modifying model architectures, unlike conventional ASR contextualization techniques (Pundak et al., 2018; Zhao et al., 2019; Jain et al., 2020; Le et al., 2021). For example, Speech LLaMA (Lakomkin et al., 2024) improved ASR performance by adding video titles and descriptions to the text prompts. Although the Speech LLaMA provides more general information along with input speech, providing individual words or phrases similarly to the conventional ASR contextualization techniques will enable LLM-based ASR systems to transcribe named entities or domain-specific terms. Specifically, some languages have many homonyms, such as Japanese and Chinese (especially when dropping tones); individual words or phrases provided as prior information help LLMs to distinguish such homonyms. Therefore, we develop an LLM-based ASR system that takes keywords as prior information.
The main contributions of this paper are that (1) we develop LLM-based ASR systems that take keywords as prior information in text prompts, and (2) experimentally show that providing keywords as prior information improves the recognition performance of rare and ambiguous words.
2 Architecture
2.1 Model Design
We adopt an in-house 100B-parameter LLM pre-trained on Japanese and English datasets from scratch, referred to as PLaMo-100B [^1] The details of this model will be published. The blog post briefly explaining this model and datasets is available at https://tech.preferred.jp/ja/blog/plamo-100b in Japanese., as a decoder, and the encoder of pre-trained Whisper large-v3222We used the checkpoint at https://huggingface.co/openai/whisper-large-v3. as an audio encoder, as shown in Figure 1. We also construct an LLM-based ASR system using Swallow-7B (Fujii et al., 2024) as a decoder that performed continual pre-training on LLaMA-2 7B (Touvron et al., 2023) on Japanese and English datasets, where Japanese was dominant.
Figure 1: Model architecture
[…]
Our 100B model does not fit in a single NVIDIA A/H100 GPU, whose memory size is 80GB. To overcome this limitation, we use QLoRA fine-tuning (Dettmers et al., 2023), where the encoder and decoder’s linear weights are quantized (Dettmers et al., 2023) and perform LoRA (Hu et al., 2022) based fine-tuning. Note that the linear adapter layer is not quantized because this is initialized with random weights rather than pre-trained weights.
2.2 Prompt Design
To contextualize the decoders, we feed additional prior information in the prompt (Lakomkin et al., 2024). We do not add special tokens to the decoder tokenizer like Whisper (Radford et al., 2022) to fix pre-trained embedding layers of the decoders. We design two prompt templates depending on the language: Japanese or English.
As a part of the prior information in the prompt, we give a language identifier inspired by Whisper’s language tag (Radford et al., 2022). Concretely, we give simple texts “Language: en” for English and {CJK}UTF8ipxm “言語:ja”, which has the same meaning in Japanese, for Japanese, respectively.
For further contextualization to transcribe rate and ambiguous words, we give keywords of the transcriptions, e.g., “Keywords: tokyo, machine learning, speech, large language model”, following Lakomkin et al. (2024). We use NA representing “not available” instead of keywords for samples not having keywords. Section 3.1 describes how to create keywords.
In summary, our input to the decoder for English datasets [^3] Appendix A gives the prompt template for Japanese datasets. is as follows:
<bos>[audio embeddings] Language: en ; Keywords: [keywords] ; Transcription: [transcription]<eos>,
where <bos> and <eos> are special tokens pre-defined by the decoder’s tokenizer, [audio embeddings] is a sequence of audio features extracted by the audio encoder and the adapter, [keywords] is a list of keywords as shown above, and [transcription] is a ground-truth audio transcription. At inference, we generate [transcription]<eos> conditioned on its prefix prompt, i.e., audio embeddings and text prior. Note that depending on the pre-training procedure of the decoder, the <eos> token in the prompt is set to be identical to the <bos> token to align the input format for packing (Brown et al., 2020), where multiple sentences are packed into a single sequence of tokens and <bos> is the delimiter between two sentences.
3 Datasets
We combined four publicly available datasets to fine-tune our models. Since our decoders and the audio encoder were pre-trained on at least Japanese and English datasets, we included an English dataset as one of four datasets to avoid catastrophic forgetting in the English domain. We evaluated ASR performance on four datasets covering Japanese and English. Tables 1 and 2 summarize the training and evaluation datasets. We briefly explain datasets in the following.
CommonVoice v8.0/16.1
We used Japanese subset of CommonVoice v8.0/16.1 (Ardila et al., 2020) [^4] https://commonvoice.mozilla.org/en/datasets. We used the original dev/test sets in CommonVoice v8.0 and v16.1 for evaluation. In addition, we used the train set of CommonVoice v16.1 for training.
LibriSpeech
As the English dataset, we used LibriSpeech (Panayotov et al., 2015), which is a commonly used benchmark dataset. We used the original train/dev/test sets in our experiments. Although dev/test sets are further divided into clean/other sets, that is, dev-clean/dev-other and test-clean/test-other sets, we will report experimental results for merged dev/test sets rather than individual clean/other sets for simplicity.
ReazonSpeech
We used a subset of ReazonSpeech v1 (Yin et al., 2023) for training, where we removed samples not included in ReazonSpeech v2. The ReazonSpeech dataset was automatically constructed by bootstrapping labeling processes using ASR models on Japanese television shows. As mentioned in the limitation of this dataset (Yin et al., 2023, Sec. 3), the original transcriptions of the television shows are inaccurate. In our experiments, we also faced a missing transcriptions issue, and we will discuss a few approaches to mitigate this issue in Section 3.2.
YODAS
We used the Japanese subset of YODAS v1 (Li et al., 2023), which is constructed from YouTube videos with a Creative Commons license. The subset has two further subsets depending on how transcriptions are generated: manual (ja000) or auto-generated (ja100). Even in the manual subset, some samples are not appropriate for ASR because their captions explain the whole video or background sounds. Thus we selected samples whose quality is estimated to be relatively high on a video-by-video basis. Concretely, we filtered out videos satisfying at least one of the following conditions: 1) character error rates (CERs) obtained by Whisper large-v3 of $40\%$ or higher and 2) ratios of alphabetic ([a-zA-Z]) characters of $50\%$ or higher. Then from the manual subset, we selected $10$ hours of high-quality videos ($87$ videos with 8 31383138\,3138 313 audio samples in total) as the test set by checking audio samples and transcriptions manually, and $10$ hours of videos randomly ($81$ videos with 8 18381838\,1838 183 audio samples in total) except for videos in the test set as a dev set. Finally, the complement set of the dev and test sets is the train set. As a result, the test set is clearer than the train/dev sets, and the dev set is clearer than the train set.
Table 1: Training dataset. In the training dataset, only YODAS is used twice. See Section 3.2 for the reason.
Name Language Keywords # samples Hours
CommonVoice v16.1 Japanese 9 61696169\,6169 616 $13.0$
LibriSpeech English 281 241281241281\,241281 241 $961.1$
ReazonSpeech Japanese 11 073 5201107352011\,073\,52011 073 520 18 867.518867.518\,867.518 867.5
YODAS Japanese $\checkmark$ 1 080 95310809531\,080\,9531 080 953 1 137.31137.31\,137.31 137.3
Total 13 703 5101370351013\,703\,51013 703 510 22 713.622713.622\,713.622 713.6
Table 2: Evaluation datasets
Name Set Language Keywords # samples Hours
CommonVoice v8 dev Japanese 4 12441244\,1244 124 $5.7$
CommonVoice v8 test Japanese 4 48344834\,4834 483 $6.5$
CommonVoice v16.1 dev Japanese 6 09460946\,0946 094 $7.9$
CommonVoice v16.1 test Japanese 6 09460946\,0946 094 $8.9$
LibriSpeech dev English 5 56755675\,5675 567 $10.5$
LibriSpeech test English 5 55955595\,5595 559 $10.7$
YODAS dev Japanese $\checkmark$ 8 18381838\,1838 183 $10.0$
YODAS test Japanese $\checkmark$ 8 31383138\,3138 313 $10.0$
3.1 Keyword Generation
We generated keywords for the YODAS datasets using another pre-trained language model because all the ASR datasets do not explicitly provide keywords for the transcriptions. We extracted keywords for each video and used them for all the pairs of audio samples and transcriptions derived from the same video since a pair of audio and transcription is a short clip of the original video in the YODAS datasets. Thus, not all keywords appear in their transcriptions. We believe this is a practical usage of keywords; a user defines keywords once for a single audio recording and reuses the keywords for all audio clips extracted from the recording rather than giving different keywords for every shorter clip.
To generate the keywords, we used Japanese Stable LM Base Gamma 7B [^5] https://huggingface.co/stabilityai/japanese-stablelm-base-gamma-7b model. Generated keywords differed depending on random seeds used for the inferences of the model. Due to this randomness, we extracted keywords with different random seeds multiple times and kept keywords generated $\nicefrac{{1}}{{3}}$ or more times.
3.2 Dataset Bias via Prompt
In our preliminary experiments, we found performance degradation due to keyword conditioning on the dataset because we only introduced the keywords to the YODAS dataset rather than multiple datasets during fine-tuning. We suspect that, as a result, the decoders can memorize the training dataset’s bias depending on whether keywords exist or not. Concretely, the ReazonSpeech dataset, which is the majority of the training dataset, tends to drop a part of the transcription text even if it exists in the speech part as the limitation in Li et al. (2023). Thus, if we give the placeholder keyword, {CJK}UTF8ipxm “キーワード:なし”, which is identical to “Keywords: na” in Japanese, to the prompts for samples from the YODAS dataset at inference, our fine-tuned models transcribe the audio as if a sample from ReazonSpeech; it drops the transcription partially. This missing transcription error increases CERs due to deletion errors as demonstrated in Section 4.7. This tendency was more severe with the PLaMo-100B-based system than the Swallow-7B’s counterpart in our preliminary experiments.
To mitigate this problem, we found approaches for fine-tuning and inference, respectively. For the fine-tuning step, we could reduce the bias by adding the YODAS dataset without keywords to the training dataset, not to condition the placeholder keywords on the ReazonSpeech dataset. For the inference step, we could reduce the degradation by using random keywords to replace the placeholder keyword. Unfortunately, the inference approach introduces additional latency, though it mitigates the performance degradation. Therefore, we used the approach for fine-tuning in our final experimental configuration. As demonstrated in experiments, we could only overcome this issue partially, which is future work.
4 Experiments
We verified our ASR systems fine-tuned on the constructed training dataset by conducting numerical experiments. As a controlled experiment to evaluate keywords-based contextualization, we also evaluated our ASR systems with slightly modified datasets and prompts; we did not feed keywords even if they were available and removed keyword text from prompts. We randomly shuffled keywords during fine-tuning and inference to make the model more robust regarding keyword ordering.
4.1 Baseline Model
We used pre-trained Whisper large-v3 downloaded from Hugging Face as the baseline system. Whisper has a typical transformer encoder-decoder architecture. It differs from our models in that the encoder and the decoder are connected through cross-attention instead of the adapter layer and input token embeddings. We explicitly specified the Whisper’s language tag for a fair comparison with our language identifier in the prompt. We used bf16 for its weights to fasten inference speed.
4.2 Pre-processing and Post-processing of Texts
Pre-processing
We applied Whisper’s text normalizers [^6] https://github.com/openai/whisper/tree/main/whisper/normalizers to the transcriptions and keywords to use consistent pre-processing with the Whisper baseline. In addition, we removed unnecessary whitespaces between non-ASCII characters for the Japanese datasets. After tokenization, we truncated keyword tokens so that the sum over lengths of prompt, keywords, and transcription becomes less than or equal to $300$ for our ASR systems.
Post-processing
We noticed that the pre-trained Whisper model transcribed characters or symbols that the text normalizers should remove. As a result, Whisper’s evaluation metric values were severely worse than our systems. For a fair comparison, we applied the text normalizers to the transcriptions generated by Whisper and our systems. This post-processing mainly boosted the performance of the Whisper model because our systems rarely generated such unintended characters and symbols. For example, CERs on the YODAS dev set of Whisper large-v3 were $16.88$ and $20.78$ with and without post-processing, respectively.
4.3 Optimization
We fine-tuned our models for only one epoch due to computational constraints. We used AdamW optimizer (Loshchilov and Hutter, 2019) with paged implementation by bitsandbytes [^7] https://github.com/TimDettmers/bitsandbytes and gradient checkpointing to reduce peak GPU memory to fine-tune PLaMo-100B-based models on even one H100 GPU. We used a linear warmup with a cosine annealing learning scheduler (Loshchilov and Hutter, 2017), whose warmup was $1\%$ iterations. We scaled the base learning rate with square root scaling depending on the number of samples in mini-batches (Hoffer et al., 2017): $\mathtt{lr}=\mathtt{base\_lr}\times\sqrt{\mathtt{bs}}$, where $\mathtt{bs}=\mathtt{mini\_batch\_size\_per\_gpu}\times\mathtt{num\_gpus}$. We minimized the cross-entropy loss over transcription tokens.
PLaMo-100B hyper-parameters
In AdamW’s hyper-parameters, we decreased $\beta_{2}=0.95$ from the default value $0.999$ by following Wortsman et al. (2023) because we observed that fine-tuning PLaMo-100B-based models is more likely to face loss spike issue [^8] Section B.1 shares other attempts to stabilize fine-tuning.. We used $\mathtt{base\_lr}=3.5\times 10^{-6}$. We used a mini-batch size of $6$ per GPU for both with and without keyword settings for pair comparison in terms of the number of gradient updates. We accelerated the training using distributed data-parallel (Li et al., 2020) with $72$ H100 GPUs.
We used $r=12$ and $\alpha=2\times r$ for QLoRA without dropout (Srivastava et al., 2014). As a result, the number of learnable weights is 218 234 880218234880218\,234\,880218 234 880 out of 103 098 001 920103098001920103\,098\,001\,920103 098 001 920 ($0.21\%$). Note that as an implementation trick of the PLaMo-100B architecture, self-attention’s three linear matrices for query/key/values (Vaswani et al., 2017) are parameterized as a single matrix, and its transformed features are split into query, key, and value representations. Not to modify this implementation, we introduced LoRA to the single-matrix rather than three query, key, and value matrices as in the native implementation. As a result, the number of LoRA weights for a set of query, key, and value of PLaMo-100B is $(3d_{k}\times r+r\times d_{\mathrm{model}})$ [^9] The terms $d_{k}$ and $d_{\mathrm{model}}$ are the same as in Vaswani et al. (2017). instead of $3(d_{k}\times r+r\times d_{\mathrm{model}})$ for a single head attention with the same dimensionality of independent key, query, and value representations. For the quantization of base weights for LoRA, we used double quantized with nf4 (Dettmers et al., 2023) implemented by bitsandbytes. During forward computation, quantized weights are de-quantized with the bf16 format.
Swallow-7B hyper-parameters
We fine-tuned Swallow-7b-plus-hf [^10] https://huggingface.co/tokyotech-llm/Swallow-7b-plus-hf as a smaller size decoder than PLaMo-100B. We only describe the different hyper-parameters from PLaMo-100B’s setting. We could use the default $\beta_{2}=0.999$ in AdamW’s hyper-parameters without frequent loss spikes. We used $\mathtt{base\_lr}=7.5\times 10^{-6}$ and mini-batch size of $64$ per GPU. We used $8$ GPUs for distributed data-parallel to perform a similar number of gradient updates to PLaMo-100B’s fine-tuning.
We independently introduced LoRA weights to query, key, and value matrices because of the implementation difference from PLaMo-100B. We used $r=16$ and $\alpha=2\times r$ for QLoRA without dropout. As a result, the number of learnable weights is 72 749 0567274905672\,749\,05672 749 056 out of 7 539 687 42475396874247\,539\,687\,4247 539 687 424 ($0.96\%$).
4.4 Transcription Generation
We transcribed audio with prompts using the greedy search for all baseline/fine-tuned models and datasets. For YODAS dev/test sets, we also transcribed audio without keywords to show the difference at inference depending on the existence of keywords. In our preliminary attempts, we allowed all models to generate $444$ tokens for transcriptions, which is the maximum number of tokens of Whisper large-v3 when specifying the language. However, we observed that this value caused too many insertion errors due to meaningless repetitions in several transcriptions, unnecessarily worsening evaluation metrics. To mitigate these undesirable insertions, we computed the max token lengths of dev sets for each dataset and tokenizer, as shown in Table 7 of Appendix B, and used $1.25$ times those lengths as the max token lengths for transcription generation for both dev and test sets.
4.5 Evaluation Metrics
As ASR performance metrics, we will report character error rates (CERs) for the Japanese datasets: CommonVoice and YODAS, and word error rates (WERs) for the English dataset: LibriSpeech. We will also report keyword error rates (KWERs) for the YODAS dev/test sets, where we selected keywords with a frequency of occurrence of 00 in train sets for evaluation from those generated for dev/test sets.
4.6 Results
Table 3: Evaluation metrics comparison. Reported metric values are Character Error Rates (CERs) except for the LibriSpeech dataset, whose metric is Word Error Rates (WERs). “KW” stands for keywords in the header. In the second and third columns, check marks $\checkmark$ represent the existence of keywords at the training and inference steps, respectively.
Model KW@Train KW@Inference CommonVoice v8.0 CommonVoice v16.1 LibriSpeech YODAS
Dev Test Dev Test Dev Test Dev Test
PLaMo $4.37$ $5.83$ $\mathbf{9.57}$ $\mathbf{13.30}$ $3.47$ $3.63$ $12.45$ $10.67$
PLaMo $\checkmark$ $\mathbf{4.24}$ $\mathbf{5.63}$ $9.70$ $13.67$ $4.16$ $4.38$ $12.93$ $11.29$
PLaMo $\checkmark$ $\checkmark$ $11.47$ $\mathbf{9.48}$
Swallow $4.99$ $6.77$ $10.22$ $13.98$ $4.22$ $4.04$ $12.72$ $11.18$
Swallow $\checkmark$ $5.05$ $7.24$ $10.32$ $13.85$ $4.39$ $4.41$ $13.05$ $11.06$
Swallow $\checkmark$ $\checkmark$ $\mathbf{11.37}$ $9.54$
Whisper $6.60$ $8.55$ $11.92$ $15.24$ $\mathbf{2.89}$ $\mathbf{2.95}$ $16.88$ $13.50$
Table 4: Keywords error rate comparison. “KW” stands for keywords in the header. In the second and third columns, check marks $\checkmark$ represent the existence of keywords at the training and inference steps, respectively. By using keywords for inference, the evaluation metric values become much lower than not using keywords.
Model KW@Train KW@Inference YODAS
Dev Test
PLaMo $63.77$ $52.74$
PLaMo $\checkmark$ $64.17$ $58.21$
PLaMo $\checkmark$ $\checkmark$ $24.29$ $\mathbf{10.45}$
Swallow $64.17$ $55.47$
Swallow $\checkmark$ $65.99$ $57.71$
Swallow $\checkmark$ $\checkmark$ $\mathbf{21.46}$ $10.95$
Whisper $71.66$ $61.69$
Table 3 shows CERs and WERs on dev/test sets. CERs for Japanese datasets improved by replacing the Whisper decoder with PLaMo-100B or Swallow-7B, probably because both models have a much larger number of parameters than the Whisper decoder, and training datasets for pre-training and fine-tuning of the models were dominated by Japanese text/audio data. In addition, compared to the case without keywords, CERs for YODAS dev/test sets improved significantly by giving keywords for training and inference with relative error rate reductions of $7.87\%$ (from $12.45\%$ to $11.47\%$ in CERs) for the dev set and $11.15\%$ (from $10.67\%$ to $9.48\%$ in CERs) for the test set for PLaMo-100B. On the other hand, compared to Whisper large-v3, the WERs of our models for LibriSpeech degraded due to the small number of English samples in the training dataset, though our models could avoid catastrophic forgetting in the English domain.
Table 4 shows KWERs for YODAS dev/test sets. As expected, we could significantly improve KWERs by giving keywords as prior information. Our model could not only improve transcriptions in terms of pronunciation but also could select correct homonyms for named entities, such as domain-specific terms and personal names, by contextualizing the decoder with keywords as shown in Table 5.
Comparing the results of PLaMo-100B with those of Swallow-7B in Table 3, although PLaMo-100B outperformed Swallow-7B for most of the evaluation sets except for the YODAS dev set, improvements in CERs and KWERs were smaller than expected from the increase in model size. One of the reasons is that the number of training epochs was insufficient for PLaMo-100B due to limitations on computational costs. Another reason is that speech recognition is easier than other tasks, such as speech translation and voice chat, and 7B models may be sufficient for speech recognition in terms of model sizes. Indeed, this tendency is consistent with Llama Team, AI @ Meta (2024, Table 31) reporting that LLaMA 3.1 8B and 70B-based multi-task speech processing systems’ WERs, where their difference was only $1.4$ at maximum. Our future work is fine-tuning our model with PLaMo-100B using more data and the number of epochs and extending to other tasks.
Table 5: Examples of recognition results in Japanese from YODAS dev/test sets improved by providing keywords as prior information. Keywords and corresponding recognized words are shown in bold (gothic), and recognition errors are marked by underlines.
{CJK}
UTF8ipxm Ground truth Without keywords With keywords {CJK}UTF8ipxg阿蘇ローズベリー香園{CJK}UTF8ipxmと {CJK}UTF8ipxg阿蘇ローズベリー{CJK}UTF8ipxmと {CJK}UTF8ipxg阿蘇ローズベリー香園{CJK}UTF8ipxmと {CJK}UTF8ipxgカラス天狗{CJK}UTF8ipxmのウルトラマン {CJK}UTF8ipxgカティング{CJK}UTF8ipxmのウルトラマン {CJK}UTF8ipxgカラス天狗{CJK}UTF8ipxmのウルトラマン {CJK}UTF8ipxmラズパイでhandbrake{CJK}UTF8ipxmいいですね {CJK}UTF8ipxmラズパイで{CJK}UTF8ipxgハンドブレーキ{CJK}UTF8ipxmいいですね {CJK}UTF8ipxmラズパイでhandbrake{CJK}UTF8ipxmいいですね {CJK}UTF8ipxg堂満直聖{CJK}UTF8ipxm26歳です {CJK}UTF8ipxg満直{CJK}UTF8ipxm26歳です {CJK}UTF8ipxg堂満直聖{CJK}UTF8ipxm26歳です {CJK}UTF8ipxg塚本秀春{CJK}UTF8ipxmさんが代表で鍵のレプリカを受け取り {CJK}UTF8ipxg塚本英治{CJK}UTF8ipxmさんが代表で鍵のレプリカを受け取り {CJK}UTF8ipxg塚本秀春{CJK}UTF8ipxmさんが代表で鍵のレプリカを受け取り
4.7 Dataset Bias Experiments
To demonstrate the issue discussed in Section 3.2, we conducted a controlled experiment using Swallow-7B models. We fine-tuned the same pre-trained Swallow-7B model, represented by cross marks ✗ in Table 6, on another training dataset that always feeds keywords for the YODAS train set; samples with keywords in the YODAS train set were shown twice to the model in the single epoch. Table 6 reports KWERs and CERs on the YODAS dev set. KWERs for transcriptions with keywords improved by using the model fine-tuned on the new train set represented by ✗, possibly by increasing the amount of training data with keywords. However, KWERs and CERs were worst when keywords were not provided to the model at inference (the second row from the bottom of Table 6), especially regarding deletions, which almost doubled. As described in Section 3.2, we suspect this is because only the YODAS train set has keywords, and other samples, mainly ReazonSpeech, do not. Thus, ReazonSpeech’s transcription error, which tends to drop a part of the transcription text, hurt the transcription of even YODAS dev/test sets by not giving keywords at inference.
Table 6: Detailed error analysis on the YODAS dev set with Swallow-7B models. The cross marks ✗ represent the use of the YODSA train set with keywords only. Recall that check marks $\checkmark$ represent the use of the YODAS train set with and without keywords, and no marks represent the use of the set without keywords only. The amount of data of the train set is the same for all conditions; each audio sample in the YODAS train set is shown to the model twice in a single epoch.
KW@Train KW@Inference KWER CER Insertions Deletions Substitutions
$64.17$ $12.72$ 6 10561056\,1056 105 6 96169616\,9616 961 10 5191051910\,51910 519
$\checkmark$ $65.99$ $13.05$ 6 42864286\,4286 428 6 93269326\,9326 932 10 8281082810\,82810 828
$\checkmark$ $\checkmark$ $21.46$ $11.37$ 5 99159915\,9915 991 6 06860686\,0686 068 9 01490149\,0149 014
$72.47$ $16.69$ 64 32643264\,3264 32 12 6211262112\,62112 621 11 8861188611\,88611 886
$\checkmark$ $18.62$ $11.65$ 6 55665566\,5566 556 6 03760376\,0376 037 8 99789978\,9978 997
Recall that we added both the YODAS dataset without and with keywords to mitigate the issue of biasing to ReazonSpeech for the training dataset with keywords for fine-tuning, which is represented by check marks $\checkmark$ at “KW@Train” in Tables 3, 4 and 6. However, comparing CERs without keywords of PLaMo-100B models in Table 3, the model fine-tuned without keywords outperformed the model fine-tuned with keywords on CommonVoice v16.1, LibriSpeech, and YODAS datasets. We suspect that the size of ReazonSpeech is still much larger than other datasets, and adding YODAS without keywords was not sufficient to completely solve this issue, resulting in performance degradation of fine-tuning with keywords compared to fine-tuning without keywords, where YODAS without keywords was used twice in a single epoch to make the train data size equivalent. We suggest diversifying the training data as much as possible if the goal is to train generalized ASR systems that can take prior texts via prompts.
5 Conclusion
We constructed an ASR system based on an in-house LLM PLaMo-100B combined with the Whisper large-v3 encoder. By contextualizing the LLM-based decoder using keywords directly as prior information in the text prompt, we showed that recognition performance for keywords that did not appear in the train set improved significantly in correcting pronunciations and selecting homonyms. In future work, training our models using more data, including other tasks and languages, is a promising direction to bring out the potential of our pre-trained LLMs.
Author Contributions
Kento Nozawa implemented a training and evaluation pipeline, conducted numerical experiments, and drafted the manuscript. Takashi Masuko constructed all datasets and extracted their keywords, implemented keywords error evaluation, and wrote the manuscript. Toru Taniguchi prototyped and supported training and evaluation pipelines, revised the manuscript, and managed this project.
Acknowledgments
This paper is based on results obtained from the project, “Research and Development Project of the Enhanced Infrastructures for Post 5G Information and Communication System” (JPNP 20017), subsidized by the New Energy and Industrial Technology Development Organization (NEDO).
We would like to thank (ex-) PFN/PFE members, especially Daiki Higurashi, Linsho Kaku, Toshiki Kataoka, Hiroaki Mikami, Shintarou Okada, Daisuke Okanohara, Shuji Suzuki, Seiya Tokui, Zijian Xu, and Shoichiro Yamaguchi, for their helpful discussions and/or implementation support; Toshihiko Yanase, Yuta Hirokawa, and Toru Komatsu for infrastructure-related support.
Appendix A Japanese Prompt
We used the following prompt for Japanese datasets:
{CJK}
UTF8ipxm
<bos>[audio embeddings] 言語:ja; キーワード:[keywords]; 書き起こし: [transcription]<eos>,
where an example of keywords is “東京、機械学習” and use “なし” if keywords are unavailable.
Appendix B Details of Experiments
Table 7 shows the max token length of dev sets using corresponding tokenizers. Note that the numbers included the special token’s length, specifically, the single end of the token.
Table 7: Max token lengths of transcriptions for dev sets.
Model CommonVoice v8.0 CommonVoice v16.1 LibriSpeech YODAS
PLaMo $51$ $54$ $122$ $61$
Swallow $52$ $54$ $132$ $62$
Whisper $60$ $69$ $117$ $71$
B.1 Attempts to Stabilize Fine-tuning
In our preliminary experiments, we sometimes encountered loss spikes during the fine-tuning that caused severe ASR performance degradation. We share our attempts to resolve these loss spikes below.
Mini-batch creation
We attempted to accelerate the training speed by using length-grouped mini-batches implemented by transformer [^11] https://github.com/huggingface/transformers/blob/v4.43.3/src/transformers/trainer_pt_utils.py#L593-L625 to make sure that each mini-batch contains similar sequence lengths. However, this causes training loss spikes especially when the length of the sequence largely changes across mini-batches. By following Morishita et al. [2017] that conducted similar experiments on neural machine translation with non-transformer models, we used random shuffling to create mini-batches instead of length grouping.
Optimizer’s hyper-parameters
Even though we used randomly shuffled mini-batches, the loss spike made the fine-tuning of PLaMo-100B models unstable. Similar to experimental configurations in Touvron et al. [2023] and Wortsman et al. [2023], reducing AdamW’s $\beta_{2}$ was effective in avoiding the loss spikes to fine-tune PLaMo-100B models.
Auxiliary loss
By following the pre-training of PLaMo-100B, we also introduced the z-loss [Chowdhery et al., 2023] with the coefficient term of $1e^{-4}$ as an auxiliary loss to stabilize fine-tuning. However, loss spikes still happened with AdamW, whose $\beta_{2}=0.999$ for PLaMo-100B, and reducing $\beta_{2}$ was helpful enough to avoid the spikes. As a result, we excluded the z-loss to avoid additional regularization.
B.2 Libraries Used in Implementation
We could conduct our experiments smoothly thanks to the following open-sourced software. All experiments were tracked by using wandb [Biewald, 2020] that GENIAC provided. In our implementations, we used transformers [Wolf et al., 2020] for encoder and decoder models’ implementations. As LoRA’s implementation, we used peft [Mangrulkar et al., 2022]. In addition, we used PyTorch [Ansel et al., 2024] for the training and evaluation pipeline and torchaudio [Hwang et al., 2023] for data-related implementation. We created our  tables by using pandas [The Pandas Development Team, 2020].